OpenStudy (anonymous):

using the sub. method.. 0.1x+0.9y=-6.8 and 0.6x-0.7y=8

OpenStudy (anonymous):

finding the solution set..

OpenStudy (anonymous):

first ditch those decimals, multiply everything by ten: 0.1x+0.9y=-6.8 and 0.6x-0.7y=8 Become: x + 9y = -68 6x - 7y =80

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Substitution: x = -68 - 9y 6(-68 - 9y) - 7y = 80 Check it, then proceed

OpenStudy (anonymous):

Multiply through 6(-68 - 9y) - 7y = 80 -408 - 54y - 7y = 80

OpenStudy (anonymous):

Combine like terms: -408 - 54y - 7y = 80 - 61y = 80 + 408 -61y = 488

OpenStudy (anonymous):

y = 488/(-61) y = -8

OpenStudy (anonymous):

And it is -8 whether you multiplied by 10 in the very first step or not. Now find y.

OpenStudy (anonymous):

i couldnot find the sub. #

OpenStudy (anonymous):

got it thanks..